# Exploring Holomorphic Retracts

### v1

@@ 1. Introduction
In particular, not every two dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the $\mathbb{B}^{N}$. We must not be amiss to note however that every one dimensional linear subspace of a bounded balanced convex domain in $\mathbb{C}^{N}$ is a retract – this follows by the Hahn–Banach theorem. It is natural to ask whether this holds if we relax the convexity condition on the domain to pseudoconvexity. The answer is in the negative, as was noted by Lempert in [Lempert-FundArt] by indicating an example to complement his passing observation therein that the existence of $2$-dimensional retracts is rather restricted and to the effect that when convexity is relaxed, even $1$-dimensional retracts can fail to exist in abundance. The only rigorous record of the just-mentioned Lempert’s example is in [Jrncki_invrnt_dst] (specifically in example 11.3.10 therein). While this relies on the usage of complex geodesics and is done only for one particular direction, we found that we can record a simpler example wherein we can begin to concretely address the question about retracts of higher dimension as well; moreover this can be achieved entirely through elementary means. In-fact, our example is as simple as the $l_{q}$-‘ball’ for $q<1$ and we record this in the following proposition. Let us also note that while the above-mentioned example of Lempert is a pseudoconvex complete Reinhardt domain much as ours, we believe recording the aforementioned $\ell^{q}$-‘ball’ is worthwhile due to the fact that the literature about retracts of dimension higher than one, is scant at-least compared to complex geodesics and one-dimensional retracts.
@@
[…]
Consequently, the only $r$-dimensional linear retracts of $D_{q}$ are those spanned by an $r$-element subset of the standard basis $\{e_{1},\ldots,e_{N}\}\subset\mathbb{C}^{N}$. Thus, the holomorphic retracts through the origin of $D_{q}$ are precisely those linear subspaces spanned by some subset of the standard basis vectors.
Next, we get back to the discussion about Vesentini’s result. Let us remark a word or two, about concrete examples of domains satisfying the $\mathbb{C}$-extremality condition in his result. Firsly, all $\ell^{q}$-balls for $1\leq q<\infty$ satisfy this condition, whereas a product of any two of them fails to, much as the polydisc. While the product operation fails to generate further examples of balls with $\mathbb{C}$-extremal property, intersection does i.e., the intersection of a pair of balls, possibly with respect to different norms, with $\mathbb{C}$-extremal boundary is again a ball (with respect to some norm) with $\mathbb{C}$-extremal boundary, as will be verified in proposition 3.1 in section 3. Vesentini’s result applies to pin-down the retracts through the center of such a ‘ball of intersection’.
Next, we may want to go beyond the class of domains in Vesentini’s result. As the most natural class of domains in several complex variables are pseudoconvex domains, it is only natural to ask if convexity in the above theorem of Vesentini can be relaxed to pseudoconvexity and the extremality condition dropped. Towards this, we have the following result wherein we assume boundedness but it must be noted that it may be replaced by hyperbolicity (allowing unboundedness), which however is not significant here, owing to an old result of Kodama in [Kodama] according to which any hyperbolic balanced domain is equivalent to a bounded one.
@@
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\longrightarrow Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin. Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a (single) holomorphic function over $D_{L}$.
@@
Owing to the assumption of being balanced, we recall that the domains as in this corollary are the much studied bounded symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV alongwith two exceptional cases, all of whose automorphism groups are well-known (cf. Hua [Hua]). This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball (as in theorem 1.7 for the $l^{p}$-ball for $p<\infty$ and theorem 1.8 for the $p=\infty$ case much earlier above), is in general quite difficult. Moreover, it must be remembered that an orthogonal projection of such $l^{p}$-balls onto a linear subspace in general is not a retract, except for the case $p=2$. As these are classical facts, the significance of linear retracts from the viewpoint of recent research work may need a remark; let us only mention that it was shown by Mok in his fairly recent work [Mok], that every linear totally geodesic submanifold through the origin of any bounded symmetric domain $\Omega$ with respect to a ${\rm Aut}(\Omega)$-invariant Kähler metric (for instance, the Bergman metric) is realizable as a retract just by orthogonal projection. As this has been dealt with deeply in [Mok], we shall proceed forth next, to look at another direction (but still related to the foregoing corollary in a sense as shall be seen), in relaxing the assumptions in Vesentini’s theorem, wherein which we retain convexity but drop the extremality condition and ask if finer information can be obtained about the retracts than the general one in the foregoing theorem 1.10. It must first be noted that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifeslty seen in the simplest case of polydiscs. This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in simple cases such as domains decomposable into a product of balls (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc. Pursuing this line of enquiry has led us to the result laid down next, which generalizes the result for polydiscs due to Heath and Suffridge, to a much larger class of domains. This theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like. We shall address this here, to the extent we can.
@@
We cannot trivially extend this theorem to the case of a product $D=B_{1}\times B_{2}\times\ldots\times B_{n}$ – where $B_{j}$’s are all bounded balanced convex domains each of whose boundaries satisfy the strict convexity condition – just by replacing one of the factors in the above theorem by a product of $(n-1)$ of these balls i.e., writing for instance $D=(B_{1}\times B_{2}\times\ldots\times B_{n-1})\times B_{n}$, appealing to the above theorem and using an inductive argument, owing to the fact that $B_{1}\times B_{2}\times\ldots\times B_{n-1}$ fails to satisfy the strict convexity condition stipulated in the above theorem as soon as $n>2$. However, such an extension is indeed possible and not difficult either; provided only that we repeat the same arguments multiple times by projecting onto the individual factors $B_{j}$ to harness the strict convexity condition satisfied by each of these factor balls of the product $D$. As the complete formulation of the end result of this, turns out to be lengthy, we shall just lay it down in section 3. Here we record an application to the union problem for a nice class of polyballs which is interesting owing to being a case with some degeneracy. Recall that the union problem comprises the following: let $M$ be a complex manifold which is the union of an increasing sequence of open subsets $M_{j}$ each of which is biholomorphic to a fixed domain $\Omega\subset\mathbb{C}^{N}$; describe $M$ in terms of $\Omega$.
@@
Next, as most balls fail to be homogeneous, the problem of characterizing retracts of bounded balanced pseudoconvex domains not passing through the origin, is all the more significant. Towards this, we report that the following result.
@@
Finally, retracts of $\Omega_{H}^{A}:=\Omega_{H}\setminus A$ where $A$ is any analytic subset of $\Omega_{H}$ are precisely given by $Z\cap\Omega_{H}$ where $Z$ is a retract of $\Omega_{H}$.
@@
If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of ×$\mathbb{C}\times D$ is either of the form ×$\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$.
@@
Let $X=\mathbb{C}\setminus A$ where $A$ is any discrete set (allowed to be empty as well), and $D$ any bounded domain in $\mathbb{C}^{N}$. Then, every retract $\widetilde{Z}$ of $X\times D$ is of the form
@@
Every non-trivial retraction map $F$ of ×$\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
@@
Then, every holomorphic retract $\widetilde{Z}$ of $X\times Y$ is either of the form
@@
where $Z$ is a retract of $Y$ and $f$ is a holomorphic map from $Z$ to $X$, or of the form $X\times Z$ where again $Z$ is a retract of $Y$.
Finally, as one concrete but ‘extreme’ example of an unbounded domain, we may simply take the domain to be $\mathbb{C}^{2}$ – indeed, this is the simplest of balanced domains of holomorphy or for that matter, the simplest of domains with non-trivial retracts and thereby too special to be remiss of us, if we do not remark about this atleast – and ask for a complete characterization of all of its holomorphic retracts (which seems unknown). Despite the apparent simplicity of the domain, this seems rather non-trivial; indeed, while graphs (in $\mathbb{C}^{2}$) of entire functions on $\mathbb{C}$ are retracts and the fact that retracts of products of planar domains in the foregoing, have literally turned out to be graphs of holomorphic maps over one of the factors, it is not the case that: every retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ is the graph of some entire function on either of its factors, as we shall show by providing a concrete example in the last section 9. On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this is attained by the following theorem.
Theorem 1.20.
If $R$ is a non-trivial polynomial retract of $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ on $\mathbb{C}^{2}$ such that $\psi(R)=\{(\zeta,0)\in\mathbb{C}^{2}\;:\;\zeta\in\mathbb{C}\}$.
Acknowledgements
This article forms the thesis work of the second author and we thank Prof. Kaushal Verma for giving us this problem of determining holomorphic retracts; as well as for helpful discussions, for which we also thank Dr. J. Jaikrishnan.

### v6

@@ 1. Introduction
It is immediate that through any given point of the polydisc, there exist retracts of all possible dimensions ($k=0,1,\ldots,N$), as its automorphism group acts transitively; a little contemplation shows that the transitive action is unnecessary to confirm the existence of such retracts. Indeed, for $1\leq p<\infty$, the $\ell_{p}$-ball fails to be homogeneous unlike the polydisc as soon as $p\neq 2$; but it is easy to derive as a corollary of theorem 1.7 that through each point of the $\ell_{p}$-ball, there exist retracts of all possible dimensions, as the retracts through the origin serve to confirm this. It is too much to as for retracts along any given set of directions because as the above proposition shows: not every two dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the ball with respect to the $l^{2}$-norm, $\mathbb{B}^{N}$. We must not be amiss to note however that every one dimensional linear subspace of a bounded balanced convex domain in $\mathbb{C}^{N}$ is a retract – this follows by the Hahn–Banach theorem. It is natural to ask whether this holds if we relax the convexity condition on the domain to pseudoconvexity. The answer is in the negative, as was noted by Lempert in [Lempert-FundArt] by indicating an example to complement his passing observation therein that the existence of $2$-dimensional retracts is rather restricted and to the effect that when convexity is relaxed, even $1$-dimensional retracts can fail to exist in abundance. The only rigorous record of the just-mentioned Lempert’s example is in [Jrncki_invrnt_dst] (specifically in example 11.3.10 therein). While this relies on the usage of complex geodesics and is done only for one particular direction, we found that we can record a simpler example wherein we can begin to concretely address the question about retracts of higher dimension as well; moreover this can be achieved entirely through elementary means. In-fact, our example is as simple as the standard $\ell^{q}$-‘ball’ for $q<1$ and we record this in the following proposition; an interesting feature of the $\ell_{q}$-‘ball’ in $\mathbb{C}^{N}$ is that apart from failing to be convex, it fails to be concave unlike its real counter-part. Let us also note that while the above-mentioned example of Lempert is a pseudoconvex complete Reinhardt domain much as ours, we believe recording the aforementioned $\ell^{q}$-‘ball’ is worthwhile due to the fact that the literature about retracts of dimension higher than one, is scant at-least compared to complex geodesics and one-dimensional retracts. In that regard, we would like to mention the very recent article of [GZ] which appeared on the arXiv, a couple of months after the earlier versions of our article. Ghosh – Zwonek attain in this article [GZ], a complete characterization of the two dimensional retracts (and thereby all retracts) of the three dimensional Lie ball and the tetrablock; this article also has noteworthy general results about retracts of Lempert domains. We now get to the special case of the $\ell^{q}$-‘ball’ for $q<1$ in arbitrary (finite) dimension, which is amongst the simplest examples of (pseudoconvex) domains which fail to be a Lempert domain (see proposition 5.8 below as well); indeed, we deal with large collections of domains that fail to be Lempert domains (however, it goes without saying that the case of Lempert domains is indeed of interest).
@@
Let $q$ be an arbitrary positive number less than one. Consider the $\ell^{q}$-‘ball’ given by $D_{q}=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;|z_{1}|^{q}+\ldots+|z_{N}|^{q}<1\}$ and the $r$-dimensional subspace $Y=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r}~{}\text{for}~{}r+1\leq i\leq N\}$ of $\mathbb{C}^{N}$, where for each fixed $i$, not all $a_{i,j}$’s are zero. Then the $r$-dimensional subspace $Y_{q}=D_{q}\cap Y$ is not a retract of $D_{q}$.
In-fact, the only $r$-dimensional linear retracts of $D_{q}$ are those spanned by an $r$-element subset of the standard basis $\{e_{1},\ldots,e_{N}\}\subset\mathbb{C}^{N}$. Thus, the holomorphic retracts through the origin of $D_{q}$ are precisely those linear subspaces spanned by some subset of the standard basis vectors.
Note that the boundaries of these domains are holomorphically extreme; indeed $\mathbb{C}$-extremal, as is the case for $\ell^{p}$-balls for $1\leq p<\infty$ as well, in which case Vesentini’s theorem reduces considerations from holomorphic retracts to linear ones. We also furnish examples (of balanced non-convex pseudoconvex domains) for the case where such extremality condition fails in the section where we detail the proof of the above theorem. Indeed, we shall show in such examples that retracts of all possible dimension exists through each point of the domain other than the origin as well. That retracts cannot also be expected in any given direction for the non-convex case, is substantiated by $D_{q}$ as in the above theorem, owing to the spectacular failure even at the origin (of $1$-dimensional retracts for instance, along any direction other than the coordinate directions). Getting back to analysing a bit further, examples arising out of $\ell^{p}$-balls, let us remark a word or two, about concrete examples of domains satisfying the $\mathbb{C}$-extremality condition in his result before we pass out of it. Firstly, while the product operation fails to generate further examples of balls with $\mathbb{C}$-extremal property, intersection does i.e., the intersection of a pair of balls, possibly with respect to different norms, with $\mathbb{C}$-extremal boundary is again a ball (with respect to some norm) with $\mathbb{C}$-extremal boundary; Vesentini’s theorem applies to pin-down the retracts through the center of such a ‘ball of intersection’. Finally and more importantly, to bring out the role of the extremality condition both in Vesentini’s theorem and allied results both in the foregoing as well as the sequel (such as theorem 1.15 involving $\mathbb{R}$-extremality), let us mention that the converse of Vesentini’s theorem is also true: if all retracts through the origin of a bounded balanced convex domain are linear, then the domain is $\mathbb{C}$-extremal. In-fact, it is enough to just have linearity of the one-dimensional retracts and furthermore, a general and local version can be formulated as follows.
@@
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $p\in\partial D$. Assume that $\partial D$ is convex near $p$ or more generally, convex at $p$ with the Minkowski functional of $D$ being continuous near $p$. Suppose the only one-dimensional retract through the origin in the direction of $p$ is the linear one, then $p$ is $\mathbb{C}$-extremal i.e., there exist non-linear retracts, as soon as $p$ fails to be a $\mathbb{C}$-extremal boundary point.
Remark 1.11.
  • (i)
    That the hypotheses that $\partial D$ is convex near $p$ implies convexity of $\partial D$ at $p$ together-with the continuity of the Minkowski functional near $p$; this will be discussed during the course of the proof. It may happen that $\partial D$ is convex at some boundary point but not in any open piece of the boundary about it; to see a simple example, consider $l^{q}$-‘balls’ for $q<1$ in $\mathbb{C}^{2}$ and examine points $p$ on the coordinate axes. This gives an example wherein convexity near $p$ fails but the weaker condition of convexity at $p$ together-with the continuity of the Minkowski functional near $p$ indeed holds. Finally here let us recall that if $p$ is a non-convex boundary point then there are no one-dimensional retracts through the origin in the direction of $p$.
  • (ii)
    More importantly, a precise description of all the non-linear retracts through the origin in a bounded balanced pseudoconvex domain along the direction of a boundary point $p$ at which $\partial D$ is convex but fails to be $\mathbb{C}$-extremal – with no assumptions about the continuity of the Minkowski functional near $p$ – is attained in proposition 6.4. We record the essence of this in the following theorem.
Theorem 1.12.
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose that $\partial D$ is convex near $p\in\partial D$ but with the boundary not being $\mathbb{C}$-extremal at $p$. Let $Z$ be an arbitrary one-dimensional retract of $D$ through the origin along the direction of $p$. Then $Z$ can be realized as the graph of a holomorphic map on the linear retract ${\rm span}(p)\cap D$, taking values in a bounded balanced pseudoconvex domain within the face $F_{\mathbb{C}}(p,\partial D)$ where $F_{\mathbb{C}}(p,\partial D)$ is the complex face of $\overline{D}$ at the boundary point $p$.
Next, we wish to go beyond the purview of Vesentini’s result. To do this by starting with the simplest step, we retain the assumption of balancedness; also, as the most natural class of domains in several complex variables are pseudoconvex domains, we ask if convexity in the above theorem of Vesentini can be relaxed to pseudoconvexity with no further assumptions about convexity as in the above theorem and extremality conditions dropped as well. Towards this, we have the following result wherein we assume boundedness. It must be noted that dropping boundedness leads to the failure of the result whereas replacing boundedness by the apparently more general condition of Kobayashi hyperbolicity is insignificant here, owing to the already mentioned result of Kodama in [Kodama], according to which any hyperbolic balanced domain is bounded.
Theorem 1.13.
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\longrightarrow Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin. Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a holomorphic map over $D_{L}$.
@@
Owing to the assumption of being balanced, we recall that the domains as in this corollary are the much studied bounded symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV alongwith two exceptional cases, all of whose automorphism groups are well-known (Hua [Hua]). This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball – as in theorem 1.7 for the $l^{p}$-ball for $p<\infty$ and theorem 1.8 for the $p=\infty$ case much earlier above; and few other relevant results as in the already mentioned very recent work [GZ] which deals with among other things, Cartan domains of type IV – is in general quite difficult. Moreover, it must be remembered that an orthogonal projection of such $l^{p}$-balls onto a linear subspace in general is not a retract, except for the case $p=2$. As these are classical facts, the significance of linear retracts from the viewpoint of recent research work may need a remark; let us only mention that it was shown by Mok in his fairly recent work [Mok], that every linear totally geodesic submanifold through the origin of any bounded symmetric domain $\Omega$ with respect to a ${\rm Aut}(\Omega)$-invariant Kähler metric (for instance, the Bergman metric) is realizable as a retract just by orthogonal projection. As this has been dealt with deeply in [Mok], we shall proceed forth next, to look at another direction (but still related to the foregoing corollary in a sense as shall be seen), in relaxing the assumptions in Vesentini’s theorem, wherein which we retain convexity but drop the extremality condition and ask if finer information can be obtained about the retracts than the general one in the foregoing theorem 1.13. It must first be noted that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifestly seen in the simplest case of polydiscs. This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in the simplest of test cases such as domains decomposable into a product of balls (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc. Pursuing this line of enquiry has led us to the result laid down next, which generalizes the result for polydiscs due to Heath and Suffridge, to a much larger class of domains and with a simpler proof following the recent work [BBMV-union]. As indicated above, this theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like. We shall address this here, to the extent we can.
@@
We cannot trivially extend this theorem to the case of a product $D=B_{1}\times B_{2}\times\ldots\times B_{n}$ – where $B_{j}$’s are all bounded balanced convex domains each of whose boundaries satisfy the strict convexity condition – just by replacing one of the factors in the above theorem by a product of $(n-1)$ of these balls i.e., writing for instance $D=(B_{1}\times B_{2}\times\ldots\times B_{n-1})\times B_{n}$, appealing to the above theorem and using an inductive argument, owing to the fact that $B_{1}\times B_{2}\times\ldots\times B_{n-1}$ fails to satisfy the strict convexity condition stipulated in the above theorem as soon as $n>2$. However, such an extension is indeed possible and not difficult either; provided only that we repeat the same arguments multiple times by projecting onto the individual factors $B_{j}$ to harness the strict convexity condition satisfied by each of these factor balls of the product $D$. As the complete formulation of the end result of this, turns out to be lengthy, we shall just lay it down in section 3. Here we record an application to the union problem for a nice class of polyballs which is interesting owing to it being a case with some degeneracy; this is based on the ideas of Fornaess – Sibony in their article [Fost]. Recall that the union problem comprises the following: let $M$ be a complex manifold which is the union of an increasing sequence of open subsets $M_{j}$ each of which is biholomorphic to a fixed domain $\Omega\subset\mathbb{C}^{N}$; describe $M$ in terms of $\Omega$.
@@
Next, as most balls fail to be homogeneous, the problem of characterizing retracts of bounded balanced pseudoconvex domains not passing through the origin, is all the more significant. Towards this, we report the following result.
@@
Finally, retracts of $\Omega_{H}^{A}:=\Omega_{H}\setminus A$ where $A$ is any analytic subset of $\Omega_{H}$ are precisely given by $Z\cap\Omega_{H}^{A}$ where $Z$ is a retract of $\Omega_{H}$.
@@
If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of ×$\mathbb{C}\times D$ is either of the form ×$\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$; and, conversely every such graph is a retract.
@@
Let $X=\mathbb{C}\setminus A$ where $A$ is any discrete set (allowed to be empty as well), and $D$ any bounded domain in $\mathbb{C}^{N}$. Then, every retract $\widetilde{Z}$ of $X\times D$ is precisely of the form
@@
Every non-trivial retraction map $F$ of ×$\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is precisely: either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
@@
Then, every holomorphic retract $\widetilde{Z}$ of $X\times Y$ is precisely: either of the form
@@
where $Z$ is a retract of $Y$ and $f$ is a holomorphic map from $Z$ to $X$ or, of the form $X\times Z$ where again $Z$ is a retract of $Y$.
Finally, as one concrete but ‘extreme’ example of an unbounded domain, we may simply take the domain to be $\mathbb{C}^{N}$ itself. While it is rather hard to precisely characterize all retracts of $\mathbb{C}^{N}$ and we shall show by actual examples that it is possible for non-hyperbolic domains to possess hyperbolic retracts, it is easy to observe a qualitative general feature common to all retracts of $\mathbb{C}^{N}$ apart from their topological triviality namely, that they are non-hyperbolic to the same (maximal) extent as $\mathbb{C}^{N}$ i.e., their Kobayashi metric vanishes identically. It is also easy to see that the one-dimensional retracts of $\mathbb{C}^{N}$ are biholomorphic to $\mathbb{C}$. For more precise results, we consider $\mathbb{C}^{2}$ – indeed, this is the simplest of balanced domains of holomorphy or for that matter, the simplest of domains with non-trivial retracts – and ask for a complete characterization of all of its holomorphic retracts (which seems unknown). Despite the apparent simplicity of the domain, this seems rather non-trivial; indeed, while graphs (in $\mathbb{C}^{2}$) of entire functions on $\mathbb{C}$ are retracts and the fact that retracts of products of planar domains in the foregoing, have literally turned out to be graphs of holomorphic maps over one of the factors, it is not the case that: every retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ is the graph of some entire function on either of its factors, as we shall show by providing a concrete example in the last section 10. On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this in essence is attained by Shpilrain and Yu in [Shpilrain--Yu] wherein it is deduced from other known results about polynomial rings and their retracts; more about this in the final section.
Theorem 1.23 (Shpilrain – Yu).
If $Z$ is a non-trivial polynomial retract of $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\Phi$ on $\mathbb{C}^{2}$ such that $\Phi(Z)=\{(\zeta,0)\in\mathbb{C}^{2}\;:\;\zeta\in\mathbb{C}\}$.
Such a ‘rectifiability’ assertion cannot in general be made for all holomorphic retracts of $\mathbb{C}^{N}$ for any $N\geq 2$, as follows by a simple analysis of the proof of the main result in the work [Forstn] due to Forstneric–Globevnik–Rosay.
Acknowledgments: We thank Gautam Bharali, J. Jaikrishnan and Kaushal Verma. The second named author acknowledges the support of the CSIR-JRF fellowship provided by the Council of Scientific and Industrial Research, India.